### v1

# Multidimensional constructs and moderated linear and nonlinear factor analysis

Abstract
Multidimensional factor models with moderations on all model parameters have so far been limited to single-factor and two-factor models. This does not align well with existing psychological measures, which are commonly intended to assess 3-5 dimensions of a latent construct. In this paper, we introduce a penalized maximum likelihood approach for multidimensional MNLFA that permits moderation of item intercepts, loadings, residual variances, factor means, variances, and correlations across three or more latent factors. Our approach incorporates ridge, lasso, and alignment penalties to stabilize estimation and detect partial measurement non-invariance while preserving model interpretability. We derive closed-form analytic gradients of the likelihood, eliminating the need for costly numerical or MCMC-based approximations, and demonstrate how this dramatically improves computational efficiency. Through simulation and application, we illustrate that penalized MNLFA recovers complex moderation patterns in multidimensional constructs and provides a scalable alternative to Bayesian implementations. We conclude by discussing the theoretical implications of penalization for measurement invariance, computational considerations, and future directions for extending the framework to categorical indicators, longitudinal data, and applied research contexts.
Keywords measurement invariance  $\cdot$ differential item functioning  $\cdot$ MNLFA  $\cdot$ Bayesian  $\cdot$ multidimensional  $\cdot$ factor analysis  $\cdot$ regularization
Introduction
A foundational component of nearly all scale development processes is investigating measurement invariance across subpopulations. Invariance of scale properties is sometimes referred to as fairness in an educational setting, as defined by the Standards (AERA, APA, NCME, 2024, p. 50) or to provide validity evidence for the consistency in interpretation of scores across groups of like respondents. Many veins of research have developed to help empirically assess the degree of validity evidence in support (or against) the conclusion that scores are comparable across groups (van Erp et al.,, 2018; van de Vijver et al.,, 2019).
One way validity evidence can be provided is by evaluating the equivalence of sets of parameters in a confirmatory factor model across two or more groups (e.g., is the entire factor loading matrix equal for males and females? Or, are all item residual variances equal over time?). Researchers have provided some excellent guidance and recommendations for scale developers to use, building on this perspective (Vandenberg and Lance,, 2000). However, these methods can be clunky when trying to identify partial invariance, which is the likely scenario where only a subset of parameters vary meaningfully across groups.
The degree of the non-invariance of subsets of parameters can be of substantive interest depending on the pattern of how parameters vary. For example, in the study of psychopathology, indicators of internalizing tend to cluster into a few types of behaviors which are treated as factors to study different indicators of psychopathology for different types of people. The pattern of how indicators of psychopathology vary across groups (e.g., sex and age) has a meaningful interpretation for the development of targeted interventions. A limitation of the traditional factor analytic approach is the difficulty in identifying these patterns due to the invariance being studied from set of parameters perspective.
An alternative perspective, often called differential item functioning, focuses on individual indicators (or items) instead of groups of parameters and has been a useful approach for identifying potentially variant indicators across groups (Clauser and Mazor,, 1998). Similarities with the multiple group factor analytic approach are discussed at length elsewhere (Thissen,, 2025). A focus on indicators has been the predominant perspective in educational testing, aiming to identify and remove items that may be biased against specific student subgroups. However, in psychological assessment, the degree of variance in parameters across groups can be of direct interest, as discussed in the study of psychopathology above. A general approach that has gained attention in recent years for the study of invariance of model parameters across groups is known as moderated nonlinear factor analysis, which is a flexible framework for incorporating a design-like matrix of group/person characteristics to be predictors of factor analysis model parameters (Bauer,, 2023).
Moderated nonlinear factor analysis has been applied in many settings to identify variance in measurement model characteristics across several background characteristics simultaneously (Bauer,, 2023). Other applications sometimes aim to provide scores on factors that are adjusted for possible partial invariance (Curran et al.,, 2014). The development of these methods has expanded into evaluating measurement invariance in up to two constructs (Bauer,, 2023). Such moderated nonlinear factor analysis models allow for the simultaneous investigation of differences in latent factor parameters alongside measurement model characteristics. Moderation effects on latent factors are sometimes referred to as the impact of background characteristics on the construct of interest. Investigating impact, while controlling for possible non-invariance of a small number of measurement model parameters, is of primary substantive interest when applying moderated nonlinear factor analysis.
The applications of moderated nonlinear factor analysis have been limited thus far to only two latent factors due to the inherent complexity of ensuring the latent variable covariance matrix is valid for all subgroups implied by the design matrix of moderating factors. The current approach for two-factor models is to induce on an unbounded continuum and then use Fisher’s z-transformation to restrict the parameter to the [-1,1] bound of correlations. Extending to more than two latent factors cannot utilize the same method using independent unbounded scores transformed to the [-1,1] scale because independent applications of Fisher’s z-transformations do not ensure the entire correlation matrix is valid. Any methods for inducing impact on individual correlations must simultaneously ensure each correlation is between [-1,1] and that the matrix of correlations is positive definite. Achieving these simultaneous constraints has been a longstanding problem in the simulation of correlation matrices (Joe,, 2006; Lewandowski et al., 2009a, ; Pourahmadi and Wang,, 2015; Archakov and Hansen,, 2021) and in covariance prediction (Pinheiro and Bates,, 1996; Pourahmadi,, 1999, 2011; Bucci et al.,, 2022). Modified conceptualizations of these methods are used in this article to address the dimensional limitation of moderated nonlinear factor analysis. We generalize moderated nonlinear factor analysis to flexibly allow moderation in the latent factor correlation parameters in models with three or more factors.
The remainder of this article is organized as follows. Next, we describe the MNLFA model in more detail for one- and two-dimensional factor models. Then, we described how MNLFA can be extended to $m$-dimensional factor models, which is the primary contribution of this article. We provide examples of how the general specification can induce a variety of correlational structures for specific subgroups. We then report on an application of the described multidimensional moderated nonlinear factor model for the study of psychopathology. We end with a discussion of the model and potential limitations and areas of future study.
Moderated (non)Linear Factor Analysis
We focus this article on latent variable models with continuous indicators, but these methods generalize to binary and Likert-type responses with some modifications. To begin, consider the measurement model for items $i=1,\ldots,I$ with observed factor indicators $y$. The single factor latent variable $\eta$ for $n=1,\ldots,N$ cases can be described as
$y_{ni}$ $=g(y^{\ast}_{ni})$ (1)
$y^{\ast}_{ni}$ $=\nu_{ni}+\lambda_{ni}\eta_{n}+\varepsilon_{ni}$ (2)
$\nu_{ni}$ $=\nu_{0i}+\mathbf{x}_{n}^{t}\bm{\delta}_{\nu i}$ (3)
$\lambda_{ni}$ $=\lambda_{0i}+\mathbf{x}_{n}^{t}\bm{\delta}_{\lambda i}$ (4)
$\varepsilon_{ni}|\mathbf{x}_{n}$ $\sim N\left(0.0,\theta_{ni}\right)$ (5)
$\theta_{ni}$ $=\theta_{0i}\exp(\mathbf{x}_{n}^{t}\bm{\delta}_{\theta i})$ (6)
$\eta_{n}|\mathbf{x}_{n}$ $\sim N\left(\alpha_{n},\phi_{n}\right)$ (7)
$\alpha_{n}$ $=\alpha_{0}+\mathbf{x}_{n}^{t}\bm{\delta}_{\alpha}$ (8)
$\phi_{n}$ $=\phi_{0}\exp{\left(\mathbf{x}_{n}^{t}\bm{\delta}_{\phi}\right)}.$ (9)
The link function $g(.)$ depends on the scale of the observed indicators and for approximately continuous variables, we can simply use the identity link function. The major difference between the above model and a traditional single-factor model is the inclusion of moderation effects on model parameters (e.g, $\nu_{ni}=\nu_{0i}+\mathbf{x}_{n}^{t}\bm{\delta}_{\nu i}$ is the moderation on item intercepts). One way to conceptualize this approach to parameterizing the factor model is to think of hierarchical linear models where the intercepts and slopes (factor loadings) are outcomes of interest (see Raudenbush and Bryk,, 2002, p. 80). Shifting to a hierarchical/multi-level perspective for these models makes a direct connection between testing for non-invariance across a wide range of possible variables with well-known methods of linear modeling.
A major benefit of connecting evaluating non-invariance to linear modeling is the flexibility that allows for unique testing of potentially nonlinear relationships between person factors and non-invariance in parameters. For example, Molenaar, (2021) illustrated how generalized additive models could be used to evaluate non-invariance. This flexibility opens the door to evaluating even more complex interrelationships between model parameters and person characteristics through interactions among characteristics. Along with the benefits, evaluating non-invariance now shares many of the same concerns that arise with respect to collinearity in the design matrix for the predictors and overall model complexity. We will return to discuss collinearity and model complexity when we discuss priors later.
The one-dimensional factor model described above is a valuable baseline for extending to more complex factor structures. The simplest extension is to a two-factor model with simple structure. In such a two-factor model, the only new wrinkle lies in evaluating impact on the factor correlation. Bauer, (2017) suggested the use of the Fisher’s z-transformation of a unconstrained parameter to handle the bounding constraint of correlations after the factor variances have been separated out.
[…]
$\phi_{n11}$ $=\phi_{011}\exp{\left(\mathbf{x}_{n}^{t}\bm{\delta}_{\phi 1}\right)}$ (11)
$\phi_{n22}$ $=\phi_{022}\exp{\left(\mathbf{x}_{n}^{t}\bm{\delta}_{\phi 2}\right)}$ (12)
$\rho_{n12}$ $=1-\frac{2}{\exp{\left[2(\gamma_{012}+\bm{\delta}_{\gamma 21}\mathbf{f})+1\right]}}$ (13)
$\phi_{n12}$ $=\rho_{n12}\sqrt{\phi_{n11}}\sqrt{\phi_{n22}}$ (14)
Fisher’s z-transformation (Eq. 11) provides a straightforward method for inducing varying factor correlations separately from differences in factor variances. However, to extend the model to three or more dimensions in the latent factor structure, Fisher’s z-transformation approach does not work to jointly ensure the positive definiteness of the latent variable covariance matrix. Next, we introduce an approach to ensuring the positive definiteness that could be used in m-dimensional latent variable space.
Multiple latent factors
Extending beyond two latent factors effectively is a complex balancing act of where in the model do you allow moderation to be estimated. A fully saturated multidimensional model has moderation in all parameters–factor loadings, item intercepts, item residual variances, factor means, factor variances, and factor correlations. However, as we discuss in more detail in the Estimation & Empirical Identification section, the fully saturated model may not be estimable in all situations due to limitations of the data and model at hand. But first, we describe the fully saturated model and the process for inducing moderation in the factor correlation matrix to illustrate the potential for these models to explore a wide range of possible sources of measurement non-invariance in different parts of the measurement model. Moderation in the item parameters is relatively straightforward and can be shown as below
$y_{ni}$ $=g\left(y^{\ast}_{ni}\right)$ (15)
$y^{\ast}_{ni}$ $=\nu_{ni}+\bm{\lambda}_{ni}\bm{\eta}_{n}+\varepsilon_{ni}$ (16)
$\nu_{ni}$ $=\nu_{0i}+\mathbf{x}_{n}^{t}\bm{\delta}_{\nu i}$ (17)
$\lambda_{nim}$ $=\lambda_{0im}+\mathbf{x}_{n}^{t}\bm{\delta}_{\lambda im}$ (18)
$\varepsilon_{ni}|\mathbf{x}_{n}$ $\sim N\left(0.0,\exp(\theta_{0i}+\mathbf{x}_{n}^{t}\bm{\delta}_{\theta i})\right)$ (19)
$\bm{\eta}_{n}|\mathbf{x}_{n}$ $\sim MVN\left(\bm{\alpha}_{n},\bm{\Phi}_{n}\right).$ (20)
The moderation in the item parameters is achieved through a model specification similar to that of the single or two-factor model. A new wrinkle involves specifying how moderation is defined in the factor loading matrix to account for possible cross-loadings. Simple structure can be assumed where each item loads onto only one factor, or any number of cross-loadings can be specified. The simplest solution is to fix all non-theorized cross-loadings to zero. A more general solution is to freely estimate a subset of loadings which we believe primarily load on different factors, and then use regularizing priors on cross-loadings. The complexity of extending to greater than two factors arises with the person-specific factor covariance matrices ($\bm{\Phi}_{n}$).
Inducing moderation in the person-specific factor covariance matrices is the core limitation of multidimensional moderated nonlinear factor analysis. This program can be restated as the more general problem of predicting several covariances simultaneously (Anderson,, 1973; Pourahmadi,, 1999). One such promising approach to inducing moderation in covariance matrices is based on a modified Cholesky decomposition that separates the variances and correlations (Pinheiro and Bates,, 1996).
Decomposing the covariance structure involves first separating the factor variances and the factor correlations. The factor variances are separated out into a diagonal matrix of the variances ($\bm{\Psi}^{D}_{n}$). The factor correlations are parameterized as a lower-triangular matrix of Cholesky factors of the correlation matrix ($\bm{\Psi}^{L}_{n}$). Together, the full covariance matrix is reconstructed using $\bm{\Psi}_{n}=\left(\bm{\Psi}^{L}_{n}\right)^{t}\left(\bm{\Psi}^{D}_{n}\right)\left(\bm{\Psi}^{L}_{n}\right)$. The value of this decomposition lies in separating the factor variances from the Cholesky decomposition of the factor correlation matrix. This serves to easily induce impact on factor variances using the exponential function to maintain positive variances and to isolate the problem of inducing impact on factor correlations restricted to the Cholesky factors.
Inducing impact, or moderation, in Cholesky factors can be accomplished using a partial correlation parameterization (Joe,, 2006; Lewandowski et al., 2009b, ) or a hypersphere parameterization (Pinheiro and Bates,, 1996; Rapisarda et al.,, 2007; Pourahmadi and Wang,, 2015). First, consider a vector of unconstrained parameters $\bm{\theta}_{n}$ with support of any real number. For $M$-factors, the number elements in $\bm{\gamma}_{n}$ is $M\times(M-1)/2=M^{\prime}$. Using the partial correlation parameterization, the vector $\bm{\gamma}_{n}$ with element $\gamma_{n12},\gamma_{n13},...,\gamma_{n1M},...,\gamma_{nM(M-1)}]$ can be used to construct the Cholesky factorization of the factor correlation matrix by first transforming to partial correlations using the hyperbolic transformation, $\theta^{\ast}=\text{tanh}\left(\gamma\right)=\frac{\exp(2\gamma)-1}{\exp(2\gamma)+1}$, resulting in the intermediary constraint of [1,1][-1,1] which partially leads to the interpretation as partial correlations. The Cholesky factorization is then built from this vector of partial correlations using the following two steps, and for the sake of simplicity, we drop the person subscript $n$.
[…]
The resulting matrix $\bm{\Phi}^{L}_{n}$ can then be constructed for each person $n$ in a flexible way to induce different correlation structures depending on the design matrix.
Secondly, using the hypersphere parameterization, the vector $\bm{\gamma}_{n}$ can be used to construct the Cholesky factorization of the factor correlation matrix by first mapping from unconstrained parameters $\bm{\gamma}_{n}$ to unique angles constrained between $0,\pi]$. Moving from the unconstrained parameters to the angles can be accomplished using a simple min-max normalization ($\theta^{\ast}=\pi\times\text{logit}(\gamma)$). These angles are transformed using $\sin$ and $\cos$ to construct the Cholesky factorization:
$\psi^{L}_{ij}$ $=\begin{cases}\hfil 0&\text{if}\ j>i\\ \hfil 1&\text{if}\ 1=j=i\\ \hfil\cos\left(\theta^{\ast}_{ij}\right)&\text{if}\ 1=j<i\\ \hfil\prod_{k=1}^{j-1}\sin\left(\theta^{\ast}_{ik}\right)&\text{if}\ 1<j=i\\ \cos\left(\theta^{\ast}_{ij}\right)\prod_{k=1}^{j-1}\sin\left(\theta^{\ast}_{ik}\right)&\text{if}\ 1<j<i\end{cases}$
A key difference between the hyperspherical parameterization and the partial correlation parameterization is the function used to convert the unconstrained parameters into the constrained space needed to construct the Cholesky factorization.
As shown above, the parameterizations are similar in form, where a vector of unconstrained parameters $\bm{\gamma}$ that are mapped to the Cholesky factorization of the correlation matrix. There is no obvious reason to prefer one parameterization over the other at this time because the inducement of impact and varying correlation structures is achieved through inducing variation in $\bm{\theta}_{n}$ by means of
$\bm{\gamma}_{n}$ $=[\gamma_{n12},\gamma_{n13},...,\gamma_{nMM^{\prime}}]$
$\gamma_{nmm^{\prime}}$ $=\gamma_{0mm^{\prime}}+\mathbf{x}_{n}^{t}\bm{\delta}_{\gamma mm^{\prime}},$
and then transforming the vector of person-specific unconstrained parameters $\bm{\theta}_{n}$. However, the interpretation of impact depends on the parameterization chosen and the element of the correlation matrix that is impacted. In the partial correlation parameterization, the impact on the correlation between factors 1 and 2 involves an intermediary step of converting from the unconstrained scale to the correlation using a hyperbolic transformation. In the hypersphere parameterization, the impact on the correlation between factors 1 and 2 has two intermediary steps: (1) converting from unconstrained scale to an angle $0,\pi]$, then (2) converting from polar angle to correlation. For interpreting the impact on other correlation coefficients, one must take into account the impact on other correlations. For example, the impact on the correlation between factors 2 and 3 depends on the impact on factors 1 and 3.
Putting the pieces of the multidimensional MNLFA model results in the model illustrated in Figure 1. The diagram illustrates the nested hierarchical nature of the model and helps break down where measurement non-invariance and impact are modeled.
Figure 1: Model specification diagram for multidimensional MNLFA. Measurement non-invariance occurs when the moderation effects ($\bm{\delta}_{\nu}$, $\bm{\delta}_{\lambda}$, $\bm{\delta}_{\theta}$) on item intercepts ($\nu$), factor loadings ($\lambda$), and item residual variances ($\theta$) are non-zero for some elements of each vector/matrix of effects; Impact on factors occurs when the moderation effects ($\bm{\delta}_{\alpha}$, $\bm{\delta}_{\phi}$, $\bm{\delta}_{\gamma}$) on factor means ($\bm{\alpha}$), factor variances ($\bm{\Psi}^{D}$), and factor correlations ($\bm{\Psi}^{L}$, Cholesky factorization) are non-zero for some elements of each vector/matrix of effects.
The choice of parameterization of the factor correlation matrix determines the function $f(.)$ used to convert the unconstrained parameters $\bm{\theta}_{n}$.
Impact of a continuous variable on correlations
The potential of this approach to inducing varying correlation structures depending on person parameters is shown in Figure 2. The figure plots the induced varying correlations across a continuous design variable $x$, where on average, the correlations are $\text{cor}(f_{1},f_{2})=0.55$, $\text{cor}(f_{1},f_{3})=0.65$, and $\text{cor}(f_{2},f_{3})=0.83$; but, as shown, the magnitude of correlation can vary quite substantially as a function of a continuous design factor. Notably, although the effect of the predictor is linear in the space of the unconstrained parameters being moderated, the effect on the induced correlation is nonlinear due to the transformation. The nonlinearity is more noticeable for the correlation between factors 2 and 3 ($\text{cor}(f_{2},f_{3})$) because impact is propagated through the impact already present for the other two correlations; that is, the impact on the correlation between factors 2 and 3 is a function of the impact on the residual correlation $\text{cor}(f_{2},f_{3})$ and the impact on the correlation between factors 1 and 3 ($\text{cor}(f_{1},f_{3})$).
[…]
Estimation & Identification
Estimating moderated nonlinear factor analyses is an active area of methodological research (Brandt et al.,, 2023, 2025; Chen and Bauer,, 2024; Gottfredson et al.,, 2019; Bauer et al.,, 2020; Kolbe et al.,, 2024). A common theme across these studies involved the use of regularization to help identify the model given the complexity of the likelihood. An approach with promise involves using Bayesian methods to approximate the posterior distribution and regularizing priors such as spike-and-slab priors for the DIF parameters (Brandt et al.,, 2023, 2025; Chen and Bauer,, 2024). The R package (Belzak,, 2023) implements Bayesian moderated nonlinear factor analyses for models with up to two latent factors relying on regularizing priors for DIF parameters (Belzak,, 2023) regDIF uses the general Bayesian MCMC package Just Another Gibbs Sampler (JAGS) to sample the marginal posterior distributions. JAGS is incredible powerful and flexible, but can be painfully slow for models with large numbers of parameters. To help overcome this limitation, we implemented the multidimensional moderated nonlinear factor analysis in Stan (REF), which is theoretically more efficient and can more easily explore the complex posterior distributions. After getting a working version of the full multidimensional MNLFA model in Stan, we ran into the same problem as with JAGS: it was painfully slow, still taking days to estimate the model with reasonable sample sizes.
The severe computational time led to exploring alternative estimation approaches and identifying where the bottleneck in computation is in estimating the model. The reason the model in Stan was taking so long is that, under the hood, Stan is approximating the gradient, the first derivative of the posterior, for each model parameter, and then moving along this approximated curve to get to the next position in the posterior surface. This gradient approximation was very computationally demanding in a general multidimensional MNLFA model, leading to the excessive slowness we observed. To overcome this complexity, we sought an estimation approach that didn’t require a numerical approximation of the gradient. We returned to old school maximum likelihood methods. Because the MNLFA model assumes a multivariate normal distribution for the responses, the likelihood function is the well-known multivariate normal, and as it turns out, the gradients for all model parameters in the MNLFA have a closed form that we can make use of when estimating the model.
Traditional methods of maximum likelihood do not incorporate regularization, but we can incorporate penalization for complexity that acts as a regularizing factor (Asparouhov and Muthén,, 2024). Instead of directly working with the log-likelihood, we can optimize a weighted composite of the log-likelihood ($L(.)$) and a penalty ($P(.)$),
$L_{w}(\theta)$ $=(1-w_{0})L(\theta)-w_{0}P(\theta),$ (21)
where $w_{0}$ is the weight given to the "prior" or the penalty term, $\theta$ is the vector of all parameters being estimated, and $L_{w}(\theta)$ represents the composite log-likelihood. The penalty term can take different forms, but the method we propose is based on a splitting the parameter vector into pieces $(\theta_{0},\theta_{1})$ where $\theta_{0}$ are the non-regularized terms (e.g., baseline item intercepts, factor loadings, etc.) and where $\theta_{1}$ contains the DIF/impact parameters. The penalty is then placed only with respect to $\theta_{1}$. A few such penalties can be defined by summing up all the squared differences among the parameters in $\theta_{1}$,
$P(\theta_{1})$ $=\sum_{\forall i}\sum_{\forall j\neq i}\frac{(\theta_{i}-\theta_{j})^{2}}{\nu},\text{the ridge penality}$ (22)
$P(\theta_{1})$ $=\sum_{\forall i}\sum_{\forall j\neq i}\frac{\sqrt{(\theta_{i}-\theta_{j})^{2}+\varepsilon}}{\nu},\text{the LASSO penality}$ (23)
$P(\theta_{1})$ $=\sum_{\forall i}\sum_{\forall j\neq i}\frac{\sqrt[4]{(\theta_{i}-\theta_{j})^{2}+\varepsilon}}{\nu},\text{the Alignment penality}$ (24)
where $\nu$ controls how severe the penalty for large differences among parameters in $\theta_{1}$. Asparouhov and Muthén, (2024) describe using the alignment penalty parameterized as $\sqrt{|\theta_{i}-\theta_{j}|+\varepsilon}$, which is approximate the same as shown above; however, using the quartic root is a smooth function and easier to differentiate and thus preferred for our purposes.
The penalized log-likelihood itself is relatively straightforward to set up to maximize using numerical methods, such as using routine optimization routines in any statistical package (e.g., optim(.) or nlminb(.) in R). However, MNLFA requires using a person-wise constructed mean and covariance matrix when evaluating the multivariate normal density function. This creates a computationally complex situation where we cannot use sufficient statistics to simplify the overall likelihood expression, as is done in lavaan. Instead must evaluate the density function for each case, or person, in our dataset after constructing the person-specific model implied mean and covariance matrix for the observed data. This complexity issue is similar to those found when trying to set up routines to estimate a multilevel structural equation model using full information maximum likelihood (see Rosseel,, 2021, for a deeper discussion on this topic). However, the tricks for efficiency described by Rosseel, (2021) may not be usable to improve the efficiency of evaluating the likelihood in MNLFA because no two persons in a dataset may have the exact same combination of moderating variables ($X$) or missing data pattern; though I certainly could be mistaken about this limitation. Due to this constraint, the only major source for efficiency gains in estimating the model lies with how the gradient of the log-likelihood is evaluated.
Gradient approximation
As noted previously, evaluating the gradient of the likelihood was the limiting factor of Stan for estimating these models. A numerical approximation of the gradient of the log-likelihood is also quite time-consuming, and the approximation increases in time needed with model size and sample size. Due to this limitation of existing methods, I derived the expressions for the gradients to alleviate the need for complex approximations of the full gradient.
Using the marginal likelihood with respect to the latent factors, the expression for the person-wise likelihood contribution is
$L_{n}=-\frac{1}{2}\ln{2\pi}-\frac{1}{2}\log|\bm{\Sigma}_{n}|-\frac{1}{2}(\mathbf{y}_{n}-\bm{\mu}_{n})^{\top}\bm{\Sigma}_{n}^{-1}(\mathbf{y}_{n}-\bm{\mu}_{n}).$ (25)
Where $\bm{\Sigma}_{n}=\bm{\Lambda}_{n}\bm{\Phi}_{n}\bm{\Lambda}_{n}^{t}+\bm{\Theta}_{n}$ and $\bm{\mu}_{n}=\bm{\nu}_{n}+\bm{\Lambda}_{n}\bm{\alpha}_{n}$. The goal now is to derive the expression for the gradient (or partial derivatives) of this likelihood contribution with respect to each possible model parameter. This step can further be broken down into the baseline values for each parameter type (e.g., baseline item intercepts and baseline factor correlations) and the moderation effects on these parameters (e.g., each $\bm{\Delta}$). This wasn’t as terrible to derive as I expected, given the well-known properties of the multivariate normal distribution.
For simplicity, I drop the subscript with respect to person $n$, but I assume that all parameters are conditional at the person-level. The gradient with respect to the baseline item intercepts $\nu$ are:
$\frac{\partial L}{\partial\bm{\nu}}=\bm{\Sigma}^{-1}(\bm{y}-\bm{\mu}).$ (26)
#’ @param ri residual of observed scores minus implied mean (y - mu)
#’ @param invSigma pre-computed invserse of implied covariance matrix
grad_nu <- function(ri, invSigma){
invSigma %*% ri
The gradient for a specific item intercept is obtained by extracting the correct row from the resulting column vector of partial derivatives. The gradient of the loglikelihood function w.r.t. to the factor loadings are a bit more involved due to the factor loadings contributing to the location and covariance matrices.
$\frac{\partial L}{\partial\bm{\Lambda}_{0}}$ $=\left(\frac{\partial\bm{\mu}}{\partial\bm{\Lambda}}\right)^{t}\Sigma^{-1}(\bm{y}-\bm{\mu})+\frac{1}{2}\operatorname{tr}\left[\left(\bm{\Sigma}^{-1}(\bm{y}-\bm{\mu})(\bm{y}-\bm{\mu})^{t}\bm{\Sigma}^{-1}-\bm{\Sigma}^{-1}\right)\frac{\partial\bm{\Sigma}}{\partial\bm{\Lambda}_{0}}\right]$ (27)
$\frac{\partial\bm{\mu}}{\partial\bm{\Lambda}_{0}}$ $=\bm{\alpha}$ (28)
$\frac{\partial\bm{\Sigma}}{\partial\bm{\Lambda}_{0}}$ $=\bm{\Phi}\bm{\Lambda}^{t}+\bm{\Lambda}\bm{\Phi}$ (29)
#’ @param i focal item
#’ @param f focal factor of item i
#’ @param ri residual of observed scores minus implied mean (y - mu)
#’ @param invSigma pre-computed invserse of implied covariance matrix
#’ @param Lambda implied factor loading matrix
#’ @param Alpha implied factor means
#’ @param Phi implied factor covariance matrix
#’ @param Qsr pre-compute quadratic form of invSigma and residual (Qsr= (invSigma) (ri) (ri^t) (InvSigma) - invSigma)
grad_lambda <- function(i, f, ri, invSigma, Lambda, Alpha, Phi, Qsr){
# 1. Dimensions
p = nrow(Lambda)
# 2. Derivative of the mean
dmu <- matrix(0,ncol=p, nrow=1)
dmu[1,i] = Alpha[f,1]
grad_mean_part = as.numeric(sum(dia(dmu %*% (invSigma %*% ri))))
# 3. Derivative of the covariance
e_i <- matrix(0, nrow=p, ncol=1)
V_f = Lambda %*% Phi[,f];
e_i[i,1] = 1;
dSigma = e_i %*% t(V_f) + V_f %*% t(e_i);
grad_cov_part = 0.5 * as.numeric(trace(Qsr %*% dSigma));
# Total gradient
grad_mean_part + grad_cov_part
where again, the gradient is the specific element of the results matrix. The gradient w.r.t. the item residual variances is
$\frac{\partial L}{\partial\bm{\Theta}_{0}}$ $=\frac{1}{2}\bm{\Theta}_{0}\left(\bm{\Sigma}^{-1}(\bm{y}-\bm{\mu})(\bm{y}-\bm{\mu})^{t}\bm{\Sigma}^{-1}-\bm{\Sigma}^{-1}\right)$ (30)
#’ @param i focal item
#’ @param Qsr pre-compute quadratic form of invSigma and residual (Qsr= (invSigma) (ri) (ri^t) (InvSigma) - invSigma)
#’ @param Theta matrix of implied residual covariances
grad_theta <- function(i, Qsr, Theta){
gx = 0.5 * Theta[i,i] * as.numeric( Qsr[i,i] )
gx
The gradient w.r.t. the factor means is
$\frac{\partial L}{\partial\bm{\alpha}_{0}}$ $=\bm{\Lambda}^{t}\bm{\Sigma}^{-1}(\bm{y}-\bm{\mu})$ (31)
$\frac{\partial\bm{\mu}_{i}}{\partial\bm{\alpha}_{0}}$ $=\bm{\Lambda}$ (32)
#’ @param ri residual of observed scores minus implied mean (y - mu)
#’ @param invSigma pre-computed invserse of implied covariance matrix
#’ @param Lambda implied factor loading matrix
grad_alpha <- function(ri, invSigma, Lambda){
gx = t(Lambda) %*% invSigma %*% ri
gx
The gradient w.r.t. to the factor variances is
$\frac{\partial L}{\partial\bm{\Phi}^{D}}$ $=\frac{1}{2}\operatorname{tr}\left[\left(\bm{\Sigma}^{-1}(\bm{y}-\bm{\mu})(\bm{y}-\bm{\mu})^{t}\bm{\Sigma}^{-1}-\bm{\Sigma}^{-1}\right)\cdot\frac{\partial\bm{\Sigma}}{\partial\bm{\Phi}^{D}}\right]$ (33)
$\frac{\partial\bm{\Sigma}}{\partial\bm{\Phi}^{D}}$ $=\bm{\Lambda}\left(\bm{\Phi}^{L}\right)^{t}\left(\frac{\partial\bm{\Phi}}{\partial\bm{\Phi}^{D}}\right)\left(\bm{\Phi}^{L}\right)\bm{\Lambda}^{t}$ (34)
$\frac{\partial\bm{\rho}}{\partial\bm{z}}$ $=1-\tanh^{2}(\bm{z})$ (35)
$\frac{\partial\bm{\Phi}}{\partial\bm{\rho}}$ $=\frac{\partial L}{\partial\bm{\rho}}\bm{\Lambda}^{t}+\bm{\Lambda}\left(\frac{\partial L}{\partial\bm{\rho}}\right)^{t}$ (36)
The gradient w.r.t. to the factor correlations is more involved due to the nested dependence of the gradients.
$\rho_{jk}$ $=\tanh(z_{jk}+\mathbf{x}^{t}\bm{\delta}_{jk})$ (37)
$\frac{\partial L}{\partial\bm{\rho}}$ $=\frac{1}{2}\operatorname{tr}\left[\left(\bm{\Sigma}^{-1}(\bm{y}-\bm{\mu})(\bm{y}-\bm{\mu})^{t}\bm{\Sigma}^{-1}-\bm{\Sigma}^{-1}\right)\cdot\frac{\partial\bm{\Sigma}}{\partial\bm{\rho}}\right]\cdot\frac{\partial\bm{\rho}}{\partial\bm{z}}$ (38)
$\frac{\partial\bm{\Sigma}}{\partial\bm{\rho}}$ $=\bm{\Lambda}\left(\frac{\partial\bm{\Phi}}{\partial\bm{\rho}}\right)\bm{\Lambda}^{t}$ (39)
$\frac{\partial\bm{\rho}}{\partial\bm{z}}$ $=1-\tanh^{2}(\bm{z})$ (40)
$\frac{\partial\bm{\Phi}}{\partial\bm{\rho}}$ $=\frac{\partial L}{\partial\bm{\rho}}\bm{\Lambda}^{t}+\bm{\Lambda}\left(\frac{\partial L}{\partial\bm{\rho}}\right)^{t}$ (41)
$\frac{\partial L}{\partial\bm{\rho}}$ $=\begin{bmatrix}1&0&\cdots&0\\ \frac{\partial L_{21}}{\partial\rho_{jk}}&\frac{\partial L_{22}}{\partial\rho_{jk}}&\cdots&0\\ \frac{\partial L_{31}}{\partial\rho_{jk}}&\frac{\partial L_{33}}{\partial\rho_{jk}}&\cdots&\vdots\\ \vdots&\vdots&\ddots&\vdots\\ \frac{\partial L_{r1}}{\partial\rho_{jk}}&\frac{\partial L_{r2}}{\partial\rho_{jk}}&\cdots&\frac{\partial L_{rs}}{\partial\rho_{jk}}\end{bmatrix},\text{where}\quad\frac{\partial L_{rs}}{\partial\rho_{jk}}:=\begin{cases}T_{jk}&\text{if }r=j,s=k,\\ -\frac{1}{l_{jj}}l_{jk}\cdot T_{jk}&\text{if }r=j,s=k,\\ \sum\limits_{l=0}^{s-1}\frac{\partial L_{rs}}{\partial\rho_{rl}}\cdot\frac{\partial L_{rl}}{\partial\rho_{jk}}&\text{if }r>j,\\ 0&\text{otherwise},\end{cases}$ (42)
where $T_{jk}$ arises from the transformation applied in constructing row $j$, and the third case expresses recursive propagation through later rows. Due to the complexity of implementing this gradient, I used a relatively computationally cheap finite differences approximation of these gradients.
#’ @param m focal element of vec(Phi_cor)
#’ @param Phi_cor_vec vector of partial correlations
#’ @param yi observed response vector for person n
#’ @param Mu model implied means of observed variables for person n
#’ @param Sigma model implied observed variable covariance matrix
#’ @param Lambda model implied factor loading matrix
#’ @param Theta model implied residual covariance matrix
#’ @param D diagonol matrix of implied factor variances
#’ @param eps (default: 1e-6) epsilon value used in finite difference approximation
finite_difference_partial_corr <- function(m, Phi_cor_vec, yi, Mu, Sigma, Lambda, Theta, D, eps=1e-6){
M = ncol(Lambda)
Phi_cor_vec_eps = Phi_cor_vec
Phi_cor_vec_eps[m] = Phi_cor_vec_eps[m] + eps
Leps = cholesky_from_partial_corrs(Phi_cor_vec_eps, M)
Sigma_eps = Lambda %*% Leps %*% D %*% t(Leps) %*% t(Lambda) + Theta
dgx = (as.numeric(dmvnrm(yi, Mu, Sigma_eps, logd = TRUE)) - as.numeric(dmvnrm(yi, Mu, Sigma, logd = TRUE)))/eps
dgx
Now for the gradients with respect to the moderation effects, $\bm{\Delta}$. The person-wise contribution to the vector of gradients with respect to the moderation effects ($\mathbf{x}^{t}$) are quite straightforward once the above mentioned gradient contributions for each baseline parameter are obtained. The form of the gradient w.r.t. for a moderating factor $\delta_{j}$ on any parameter $\theta$ (not to be confused with the item residual variance because here, $\theta$ is a placeholder for any parameter) is thus
$\frac{\partial L}{\partial\bm{\delta}_{\theta}}=x_{j}\cdot\frac{\partial L}{\partial\theta}.$ (43)
So the gradient with respect to design variable $x_{j}$ on item intercept $i$ is therefore
$\frac{\partial L}{\partial\delta_{\nu_{i}j}}=x_{j}\cdot\frac{\partial L}{\partial\nu_{i}}$ (44)
The other gradient contributions are found similarly. Another important consideration is the approximation of the hessian for approximation of standard errors.
Hessian approximation
The Hessian, or observed information matrix, of the parameter vector, $\bm{\theta}$, is a necessary contribution to approximate the uncertainty in the estimate parameters. We recommend using a sandwich-like estimator for the standard errors using the inverse of the observed information matrix and the outer-product of the gradient (Freedman,, 2006).
Discussion
The penalized maximum likelihood framework for multidimensional moderated nonlinear factor analysis (MNLFA) developed in this paper represents an important step toward addressing the long-standing limitations of existing methods for studying measurement invariance. Traditional applications of MNLFA have been restricted to single- or two-factor models, primarily because of the mathematical and computational difficulty of ensuring valid correlation structures when moderation is introduced at the factor covariance level. By leveraging penalization strategies alongside analytic gradients of the likelihood, our framework allows estimation of models with three or more latent factors in a way that balances flexibility with computational feasibility.
This work builds directly on the original MNLFA framework introduced by Curran and colleagues, who developed a moderated nonlinear factor model for integrative data analysis (IDA) in pooled longitudinal and developmental datasets (Curran et al.,, 2014). That seminal work highlighted the potential of moderated factor models to accommodate construct non-invariance across heterogeneous subpopulations, but also revealed the challenges of extending beyond two factors. Subsequent research has expanded the methodological foundation of MNLFA, particularly through the adoption of Bayesian approaches. For instance, Brandt et al., (2023, 2025) proposed Bayesian penalty methods using shrinkage priors (e.g., spike-and-slab, lasso-type, and horseshoe priors) to regularize moderation effects and detect differential item functioning (DIF) across multiple moderators. These Bayesian approaches are powerful, especially for simultaneously handling many potential sources of non-invariance, but they are often computationally demanding and require expertise in specifying priors and interpreting posterior distributions.
Our contribution complements this Bayesian line of work by providing a likelihood-based analogue that is both familiar to applied researchers and computationally efficient. In contrast to simulation-based Bayesian estimation, the penalized maximum likelihood framework exploits closed-form gradients of the multivariate normal likelihood with respect to all model parameters, including factor loadings, intercepts, residual variances, factor means, and moderated correlations. This analytic treatment avoids the computational bottleneck of gradient approximation that plagues Bayesian software such as JAGS or Stan. As a result, penalized maximum likelihood estimation is considerably faster, making multidimensional MNLFA accessible for large-scale applications where Bayesian estimation might be prohibitively slow.
The idea of incorporating penalties into factor analytic frameworks is not new, and our work is situated within a growing literature on penalized likelihood methods in psychometrics and structural equation modeling. Finch (2018), for example, demonstrated via Monte Carlo simulations that penalized likelihood estimation improves the detection of measurement non-invariance by controlling Type I error while maintaining or improving power relative to traditional multi-group CFA tests. Geminiani et al., (2021) extended penalized factor analysis to both single- and multi-group settings, introducing a general framework for sparsity-inducing penalties in confirmatory models. Hirose and Yamamoto, (2012) and Choi, (2010) developed related non-concave penalized likelihood methods for inducing sparsity in factor loadings, showing how penalization can directly enhance interpretability. Together, this body of work demonstrates that penalties are not only computationally stabilizing but also substantively meaningful, enabling more parsimonious and interpretable representations of latent constructs. Our alignment, ridge, and lasso penalties extend these ideas to the moderation context, offering applied researchers flexible options for calibrating the degree of invariance across model parameters.
Beyond theoretical contribution, penalized MNLFA has significant practical implications. Many substantive domains—including psychopathology, education, and public health—routinely study constructs that are multidimensional, often spanning three to five latent factors. Applied researchers are rarely interested in strict invariance across all measurement parameters, since such invariance is both unrealistic and unnecessary. Instead, they seek to identify where parameters differ across groups or person-level moderators, and to interpret the substantive meaning of such differences. Penalization provides a principled way to detect partial invariance without overfitting or producing unstable estimates. The alignment penalty, for example, encourages approximate invariance across parameters while still allowing localized deviations, echoing the goals of alignment optimization methods in multiple-group CFA (Asparouhov and Muthén,, 2024).
Despite these advantages, challenges remain. A primary computational bottleneck arises from the need to construct person-specific implied means and covariance matrices in the likelihood evaluation. Because moderation enters directly into item and factor parameters, sufficient statistics cannot be used to simplify the likelihood, as is possible in conventional CFA software such as lavaan. Instead, the density function must be evaluated separately for each individual, limiting scalability. While our derivations of analytic gradients alleviate much of this burden, further efficiency gains may come from algorithmic innovations such as caching strategies, low-rank approximations, or GPU acceleration. This echoes concerns raised by Lüdtke et al., (2021) , who found that small-sample confirmatory factor analysis often struggles with convergence under MCMC, but can benefit from penalized ML approaches due to their relative stability.
Another area for development is the design of penalty functions themselves. While we focus here on ridge, lasso, and alignment penalties, other formulations may offer distinct advantages. Elastic net penalties, for example, combine the strengths of ridge and lasso by balancing smooth shrinkage with sparsity induction. Adaptive penalties can weight moderation effects differently, improving bias-variance tradeoffs. Group penalties could be used to enforce structure across related sets of moderation effects (e.g., all loadings associated with a particular factor). Systematic simulation studies will be essential to compare the performance of these alternatives under varying conditions of dimensionality, sample size, and degree of true non-invariance.
Extending penalized MNLFA beyond continuous indicators also represents a promising direction. Many applied settings involve binary or ordinal indicators, longitudinal repeated measures, or mixtures of latent subgroups. While the general penalized likelihood framework extends naturally to such cases, additional derivations are required for closed-form gradients in non-Gaussian models. Software implementations are another key frontier. Current packages such as aMNLFA (Gottfredson et al.,, 2019) and mnlfa (Robitzsch,, 2023) have increased accessibility of MNLFA by automating score generation and incorporating penalization for limited-factor models. Developing similar tools for multidimensional penalized MNLFA will be critical for broader adoption. Integration into widely used SEM platforms or stand-alone R packages could help disseminate these methods to applied researchers who might otherwise be deterred by computational or coding complexity.
In conclusion, the penalized maximum likelihood framework presented here advances the methodological toolkit for evaluating measurement invariance in multidimensional constructs. It extends the seminal contributions of Curran et al., (2014) and complements recent Bayesian penalty approaches (Brandt et al.,, 2025) by offering a frequentist alternative that is both interpretable and computationally efficient. By combining analytic gradients with penalization, our approach makes multidimensional MNLFA practically estimable and theoretically robust. Challenges remain in scalability, penalty design, and software implementation, but the path forward is clear: continued development, comparative validation, and applied demonstration will help realize the full potential of penalized MNLFA for promoting fairness, validity, and precision in measurement across diverse populations.
Acknowledgments
Many of the computations in this paper were run on the FASRC Cannon cluster supported by the FAS Division of Science Research Computing Group at Harvard University.